# The 1/3 Geometric Constant: Scale Invariance and the Origin of ’Missing Energy’ in 3D Quantum Fragmentation

### v2

@@ V Principal Quantum Number Effects and Multi-Peak Phenomena
Beyond the ground state ($n=1$), higher principal quantum numbers ($n>1$) introduce additional radial nodes and qualitative changes to the topology of the energy distribution $P(E)$. Figure 6 illustrates the computed spectra for $n=1,2,$ and $3$, maintaining fixed parameters of $Q=1$ and $\zeta=1$ with $l=0$.
V.1 Emergence of the Secondary High-Energy Peak
@@
V.2 Experimental Context: $H_{2}^{+}$ Ionization and Higher Shells
Despite the intensity of these secondary features in pure excited states, their contribution to a total experimental spectrum is typically secondary to the $n=1$ signal. In most physical systems, the population of excited states remains a small fraction relative to the ground state. Consequently, these extra peaks manifest as a “broadening” or a high-energy structured tail in the final observed spectrum.
This phenomenon provides a rigorous theoretical basis for observations in molecular dissociation. Specifically, in our forthcoming work regarding $H_{2}^{+}$ circular ionization, we demonstrate that the population of higher $n$ states correlates directly with the appearance of extra peaks in the proton Kinetic Energy Release (KER) spectrum. The consistency between our Slater-basis model and full-dimensional $H_{2}^{+}$ simulations suggests that these high-energy signatures are a universal footprint of excited-state participation in Coulomb explosions.
VI The Geometric Origin of Missing Energy: A Cautionary Note for Particle and Nuclear Physics
The persistence of the $1/3$ ratio in our scaling analysis ($M=\zeta$) suggests a fundamental re-interpretation of the ”missing energy” problem in subatomic physics. Historically, the continuous spectrum of beta decay—characterized by an average electron energy $\langle E\rangle\approx 1/3E_{max}$—was the primary catalyst for the postulation of the neutrino Ellis and Wooster (1927). Under the classical assumption, energy conservation could only be satisfied if a non-interacting particle carried away the remaining two-thirds of the energy.
However, our derivation of the $R_{E}\approx 0.33$ constant at a unit charge $Q=1$ demonstrates that this energy imbalance is an intrinsic property of the $4\pi r^{2}$ volume weighting inherent to three-dimensional quantum geometry. In this framework, the energy is not ”missing” in the sense of being evacuated from the system by an auxiliary particle. Instead, it is sequestered in the high-energy tail of the initial state’s 3D radial distribution. Because the experimental detection mode focuses on the peak (the most probable energy), it is geometrically blind to the majority of the integrated energy budget stored in the high-gradient components of the wavefunction ($\nabla^{2}\psi$).
We propose that the $R_{E}<0.5$ bound and the $1/3$ ground-state constant be treated as a ”geometric baseline” for energy calorimetry. If the $1/3$ ratio is a topological artifact of 3D space, then the ”missing” energy observed in nuclear transitions may represent the recovered high-momentum components of the primary fragment rather than a separate particle species. Consequently, we caution that invoking new degrees of freedom to satisfy energy conservation may be premature without first accounting for the $r^{2}$ suppression effect documented in this work. This perspective shifts the focus from looking for ”missing particles” to correcting the ”geometric perspective” of the detector.
VII Conclusion
In this work, we have identified a fundamental geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum systems. By analyzing the fragmentation of Slater-type orbitals—the physical standard for localized quantum states—we demonstrate that the ratio of the peak detected energy to the integrated mean, $R_{E}=E_{\text{peak}}/\langle E\rangle$, is strictly bounded below $0.5$. This discrepancy is not a dynamical error but a topological consequence of the $4\pi r^{2}$ volume element, which inherently sequesters a majority of the interaction energy in the high-gradient tails of the 3D wavefunction.
Our parametric analysis reveals two critical benchmarks for physics. First, the $0.5$ ratio acts as a universal ”geometric baseline” for 3D quantum simulations; since most complex wavefunctions are expanded in a basis of Slater-type orbitals, this bound represents a fundamental limit on what a peak-centric measurement can capture from a localized system. Second, in the high-localization limit characteristic of subatomic scales ($M\propto\zeta$), the ratio converges to a scale-invariant constant of $\approx 0.33$ for ground-state ($n=1$) transitions.
The precise agreement between our $1/3$ geometric constant and the historical ”one-third rule” of beta-decay energy distributions offers a provocative re-interpretation of the ”missing energy” problem. We suggest that the energy imbalance which necessitated the neutrino hypothesis may be, at least in part, a manifestation of the geometric volume element’s effect on energy calorimetry. By identifying the most probable energy (the peak) rather than the statistical mean, experimental observations are geometrically predisposed to underestimate the total energy budget by a factor of approximately three.
Ultimately, this work cautions that the invocation of new physical degrees of freedom to satisfy energy conservation may be premature without first accounting for the $r^{2}$ suppression effect. This geometric correction provides a necessary baseline for the accurate reconstruction of initial-state energies across atomic, nuclear, and high-energy physics, shifting the focus from the search for undetected mass to a more rigorous account of 3D quantum geometry.

### v3

@@ VI Principal Quantum Number Effects and Multi-Peak Phenomena
Beyond the ground state ($n=1$), higher principal quantum numbers ($n>1$) introduce additional radial nodes and qualitative changes to the topology of the energy distribution $P(E)$. Figure 5 illustrates the computed spectra for $n=1,2,$ and $3$, maintaining fixed parameters of $Q=1$ and $\zeta=1$ with $M=1$.
@@
VII Evidence Across Atomic and Subatomic Scales
VII.1 Empirical Validation: The $H_{2}^{+}$ Energy Discrepancy
Figure 6: Proton KER spectra for $H_{2}^{+}$ dissociative ionization. Dashed and solid lines represent 1D (taken from Ref. Yue and Madsen (2013)) and 3D computations, respectively. Triangle markers denote experimental data taken from Ref. Pavičić et al. (2005). The black curves represent ground-state ($n=1$) contributions, while red curves include contributions from populated excited states.
The physical validity of the 0.33 geometric landmark is most strikingly evidenced by the long-standing energy discrepancy in attosecond $H_{2}^{+}$ fragmentation. As shown in Fig. 6, traditional 1D semiclassical models Yue and Madsen (2013) predict a KER peak between 10101313 eV, significantly overestimating the experimental observation. In contrast, our 3D simulations (solid black line), performed via the tRecX quantum framework, yield a primary peak near $3$ eV, in excellent agreement with experimental calorimetry.
Our model replicates this $\approx 7$ eV shift without requiring external field effects or complex multi-photon dynamics. Although the $H_{2}^{+}$ system lacks perfect central symmetry and involves fragments with significant mass ($M_{p}\approx 1836$ a.u.), the geometric constraint remains the dominant physical driver. Because the initial $1\sigma_{g}$ state is highly localized, the $4\pi r^{2}$ volume weighting effectively sequesters the majority of the potential energy in the high-momentum radial tail. Consequently, the most probable detected energy ($E_{\text{peak}}$) represents only one-third of the integrated mean energy ($\langle E\rangle\approx 11$ eV), confirming that the “missing energy” in molecular physics is a direct manifestation of 3D quantum geometry.
Furthermore, when excited states ($n>1$) are involved, our 3D model identifies a secondary high-energy feature between 4455 eV (red curves in Fig. 6). This secondary peak, which is entirely absent in 1D approximations, serves as a signature of the radial nodal structures in the initial state. While excited-state populations are typically secondary to the $n=1$ signal, their presence explains the characteristic broadening and high-energy structured tails observed in experimental KER spectra.
VII.2 The Geometric Origin of Missing Energy in Subatomic Physics
Figure 7: Comparison of the historical beta decay spectrum of Radium E (210Bi), adapted from Ref. Ellis and Wooster (1927) (triangle markers), and our theoretical $P(E)$ distribution with $\alpha=0.16$ (red solid line). The spectrum exhibits the characteristic 1/3 ratio between mean and endpoint energy, consistent with our predicted geometric landmark.
The identification of the $M=\alpha\zeta/Q$ scaling law provides a provocative new lens for interpreting energy anomalies in nuclear physics. Historically, the continuous electron spectrum observed in beta decay—characterized by an average energy $\langle E\rangle\approx 1/3E_{\text{max}}$—was the primary catalyst for the neutrino hypothesis Ellis and Wooster (1927). Notably, the morphology of the beta spectrum (Fig. 7) is remarkably well-replicated by our model in the low-to-mid energy range ($0.5E_{\text{max}}$) using a coupling constant of $\alpha=0.16$. This value corresponds to a high-localization ($\zeta$) regime, consistent with the high atomic number ($Z$) of Radium.
The divergence of the fit at higher energies is expected, as this region enters the relativistic regime where the non-relativistic Laplacian used in our current $P(E)$ derivation reaches its limit of applicability. In this regime, the operator overestimates the energy—driven, causing the fit to curve above the Beta Decay baseline, as illustrated in the figure. Furthermore, the nuclear $Q$-value imposes a hard phase-space cutoff at the endpoint energy that truncates the theoretical high-energy tail. Despite these high-energy corrections, our results suggest that the $R_{E}\approx 0.33$ ratio represents a fundamental geometric baseline for energy detection in localized 3D systems. This geometric constraint potentially offers a unified explanation for perceived missing energy across diverse regimes, from Auger-induced Coulomb explosions to the quenching factors observed in dark matter direct detection Levin and others (1991); Aprile and others (XENON Collaboration) (2023).
VII.3 Unified Mechanism of Spectral Broadening and Orbital Contraction
The spectral distribution $P(E)$ is critically governed by the localization parameter $\zeta$. As illustrated in our parametric study (Fig. 1), an increase in $\zeta$ induces a rapid broadening of the high-energy tail. In high-$Z$ nuclei, the proximity of inner-shell electrons effectively scales the $\zeta$ parameter, leading to the spectral broadening observed in traditional nuclear studies Wilkinson (1990).
For superheavy nuclei ($Z>100$), the extreme localization (high $\zeta$) increases the probability of detecting ultra-high energy fragments. This phenomenon aligns with experimental observations of relativistic orbital contraction Henning (2012). While currently interpreted via Fermi theory Pershina (2016), our model suggests these relativistic shifts are consistent with the geometric pressure inherent to high-$\zeta$ localized states, providing a scale-invariant link between atomic and relativistic nuclear dynamics.
VIII Conclusion
In this work, we have demonstrated that the persistent discrepancy between theoretical energy budgets and experimental peaks in quantum fragmentation is not merely a technical limitation, but a fundamental topological artifact of three-dimensional wavefunctions. By utilizing localized Slater-type orbitals and full-dimensional TDSE benchmarks, we have established that the $4\pi r^{2}$ volume element inherently masks a significant portion of a system’s kinetic energy, imposing a universal peak-to-mean energy bound of $R_{E}<0.5$.
The discovery of the scale-invariant law, $M=\alpha\zeta/Q$, provides a unified framework for energy reconstruction across twelve orders of magnitude. We have shown that the observed 1/3 ratio in $H_{2}^{+}$ dissociative ionization is a direct manifestation of this geometric constraint, replicated here without the need for complex multi-photon dynamics or external field corrections. Most significantly, the alignment between this “Geometric Zero-Point” and the historical average-to-endpoint discrepancy in beta decay suggests that 0.33 is a universal landmark of 3D detection.
Our findings offer a cautionary note for the interpretation of “missing energy” in particle and nuclear physics. While the neutrino and other non-interacting particles are essential components of modern theory, our results reveal that the baseline for energy detection in localized subatomic states is inherently skewed by the topology of space itself. Because the $1/3$ ratio persists—and can even decay toward zero—as particles become more massive and localized, this geometric masking must be decoupled from signatures of invisible flux. This perspective shifts the focus of future calorimetry from the search for undetected mass to the correction of the intrinsic geometric filters of the three-dimensional quantum world.
Appendix A Fiting curve of $P(E)$
To identify a universal analytical form for the fragment energy spectrum, we performed a regression analysis on the energy distribution P(E) using several statistical models. Fig. 8 illustrates the results for a ground-state hydrogenic orbital ($\zeta=1,Q=1$) and a more localized, higher-charge configuration ($\zeta=5.2,Q=1.6$).
We compared the computed data (blue) against three candidate distributions: a radial-type distribution (r-fit), a Gamma distribution, and a $\chi^{2}$ distribution. As shown in Fig. 2, the r-fit (green) and the Gamma distribution (cyan) produce the highest fidelity, with a residual error of less than 2%. The near-perfect overlap between these two functions confirms that the energy distribution essentially obeys a radial distribution law.